# Tele-rehabilitation with online skill transfer and adaptation in $\mathbb{R}^{3}\times\mathit{S}^{3}$

### v1

@@ III-A Teleoperation Control Architecture
The above architecture demands a reliable communication channel to transfer the pose of both robots bilaterally. Thus, a position-position (PP) structure of teleoperation is implemented here to ensure the synchronization of the TR and PR.
@@ III-B Motion generation
decides the shape of DMP. $c_{i}$ represents the center of radial basis function (RBF), $h$ denotes the kernel width, and $r$ is the amplitude modulation parameter which is set to $1$. NN is the number of RBF. Each vector $\bm{w}_{i}\in\mathbb{R}^{6}$ of the weights matrix $\bm{w}\in\mathbb{R}^{6\times N}$ is updated with the Recursive Least Square (RLS) algorithm [24]. In practice, different number of RBF functions and width for translational and rotational components can be chosen. The forgetting factor $\lambda_{fg}$ in RLS decides how much recent data is used for updating the weights[25]. The larger the forgetting factor, the more recent the data is used, but the slower the convergence.
@@ III-C Autonomy Allocation
where $\lambda_{err}$ is the hyperparameter serving as tolerance of learning error. The parameter $I_{s}$ is a value indicates the confidence of the learned skill. When the error decreases below $\lambda_{err}$, the term $(1-I_{s})$ becomes a positive number, which will gradually drive $\mu$ to $1$ and lower the learning rate in Equation 23 until zero. When the error is larger than $\lambda_{err}$, the $(1-I_{s})$ becomes negative and force $\mu$ to zero, which increases the learning rate to one.
@@ IV-A Robot Experiment
The experiment encompasses the demonstration of five periodic rehabilitation skills denoted as $R_{1},\cdots,R_{5}$ with translational $XYZ$ in Euclidean space and orientational motion expressed by a quaternion. Specifically, the skills contain,
@@ IV-B Sensitivity Analysis
A trade-off emerges between DMP weight convergence, measured by the weight standard deviation, and the steady-state error. We vary the basis-function width as $h\in\{1,3,8,31,100\}$, corresponding to overlaps with approximately $6$, 33, 22, 11, and $0.5$ neighboring basis functions, respectively. As shown in Table I, larger widths (i.e., less overlap) improve weight convergence but increase the steady-state error. Balancing these effects, we select $h=31$.

### v2

@@ III-A Teleoperation Control Architecture
The above architecture demands a reliable communication channel to transfer the pose of both robots bilaterally. Thus, a position-position ($P$) structure of teleoperation is implemented here to ensure the synchronization of the TR and PR.
@@ III-B Motion generation
decides the shape of DMP. $c_{i}$ represents the center of radial basis function (RBF), $h$ denotes the kernel width, and $r$ is the amplitude modulation parameter which is set to 11. NN is the number of RBF. Each vector $\bm{w}_{i}\in\mathbb{R}^{6}$ of the weights matrix $\bm{w}\in\mathbb{R}^{6\times N}$ is updated with the Recursive Least Square (RLS) algorithm [24]. In practice, different number of RBF functions and width for translational and rotational components can be chosen. The forgetting factor $\lambda_{fg}$ in RLS decides how much recent data is used for updating the weights[25]. The larger the forgetting factor, the more recent the data is used, but the slower the convergence.
@@ III-C Autonomy Allocation
where $\lambda_{err}$ is the hyperparameter serving as tolerance of learning error. The parameter $I_{s}$ is a value indicates the confidence of the learned skill. When the error decreases below $\lambda_{err}$, the term $(1-I_{s})$ becomes a positive number, which will gradually drive $\mu$ to 11 and lower the learning rate in Equation 23 until zero. When the error is larger than $\lambda_{err}$, the $(1-I_{s})$ becomes negative and force $\mu$ to zero, which increases the learning rate to one.
@@ IV-A Robot Experiment
The experiment encompasses the demonstration of five periodic rehabilitation skills denoted as $R_{1},\cdots,R_{5}$ with translational XYZXYZ in Euclidean space and orientational motion expressed by a quaternion. Specifically, the skills contain,
@@ IV-B Sensitivity Analysis
A trade-off emerges between DMP weight convergence, measured by the weight standard deviation, and the steady-state error. We vary the basis-function width as $h\in\{1,3,8,31,100\}$, corresponding to overlaps with approximately 66, 33, 22, 11, and $0.5$ neighboring basis functions, respectively. As shown in Table I, larger widths (i.e., less overlap) improve weight convergence but increase the steady-state error. Balancing these effects, we select $h=31$.